OpenStudy (anonymous):

if the slope of line p is 1/5 and the slope of line q is -5, then what is true abou the lines p and q (assume lines p and q are coplanar)?

OpenStudy (anonymous):

they are parallel they are perpendicular they are neither parallel nor perpendicular they are skew lines

OpenStudy (whpalmer4):

parallel lines have the same slope perpendicular lines have slopes whose product is -1

OpenStudy (anonymous):

so is it A?

OpenStudy (whpalmer4):

isn't A "they are parallel"? Doesn't the problem say that one line has slope 1/5 and the other line has slope -5? Didn't I say parallel lines have the same slope? Are 1/5 and -5 the same?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-5=1/5

OpenStudy (whpalmer4):

-5 does not equal 1/5!!!

OpenStudy (anonymous):

oh ok so its not A either then what is it

OpenStudy (whpalmer4):

are we writing your name on the paper, or mine? :-)

OpenStudy (whpalmer4):

what is -5 * 1/5?

OpenStudy (anonymous):

1?

OpenStudy (whpalmer4):

5 * 1/5 = 1 -5 * 1/5 = ?

OpenStudy (whpalmer4):

a negative number * a positive number gives you a negative number a negative number * a negative number gives you a positive number a positive number * a positive number gives you a positive number

OpenStudy (whpalmer4):

-5 * 1/5 = -1 as I mentioned earlier, if two lines' slopes multiplied together give you -1, the lines are perpendicular

OpenStudy (anonymous):

oh so b?

OpenStudy (whpalmer4):

if b is the answer that says they are perpendicular, then yes

OpenStudy (anonymous):

thnx and sorry im so stupid XD

OpenStudy (whpalmer4):

not stupid, just more room for improvement :-)